# Correlation between $\mathrm{H}\rm{\alpha}$ emitters and their cosmic web environment at $z\sim 1$

### v1

@@ 3.3 Cosmic web classification
The top panel of Fig. 5 displays the environmental mark fractions of ELG parent halos as a function of the ELG total $\mathrm{H}\rm{\alpha}$ luminosity. In each luminosity bin, we computed the fraction of galaxies in every classification, such that the fractions sum to 1 by definition. The threshold values $\lambda_{\text{th,V}}$ and $\lambda_{\text{th,T}}$ adopted here ensure that the V- and T-web classifications are in reasonably good agreement with those predicted by DisPerSE for $\mathrm{H}\rm{\alpha}$ luminosities greater than $10^{41}\ {\rm erg\,s^{-1}}$. Observe also that the fraction of ELGs in filaments is nearly constant and equal to $\approx 40\%$ across four orders of magnitude in luminosity (this is more obviously true for V-web and DisPerSE than for T-web), unlike the fraction of galaxies in sheets. At $L_{\mathrm{H}\rm{\alpha}}>10^{43}\ {\rm erg\,s^{-1}}$, nearly all galaxies are in knots. Finally, note that ELGs with parent halo mass smaller than $2\times 10^{11}\ {\rm{\it h}}^{-1}M_{\odot}$ are not included in the classification reported in Fig. 5 due to the limited resolution of the grid. This excludes only $\approx 18\%$ ($2\%$) of ELGs with $L_{\mathrm{H}\rm{\alpha}}>10^{39}\ {\rm erg\,s^{-1}}$ ($L_{\mathrm{H}\rm{\alpha}}>10^{42}\ {\rm erg\,s^{-1}}$).
3.4 Environmental bias
@@
The peak-background split argument relates the magnitude of the large scale bias $b_{1}$ to (minus) the slope of the halo mass function (Kaiser, 1984; Bardeen et al., 1986; Cole & Kaiser, 1989; Mo & White, 1996; Sheth & Tormen, 1999). In light of this, Fig. 4 suggests that, at fixed halo mass, halos in sheets and filaments should be significantly more clustered than those residing in knots.
@@
To illustrate this point, the top panel of Fig. 6 shows the distribution of the anisotropy parameters $\alpha_{\rm V}$ and $\alpha_{\rm T}$ for halos with mass $M\geq 5\times 10^{11}\ M_{\odot}/h$. The anisotropy parameter thus is a good discriminator of the environment. Furthermore, the largest $\alpha_{\rm V}$ and $\alpha_{\rm T}$ values are exclusively associated with the rare halos residing in sheets. The bottom panel of Fig. 6 displays the corresponding bias transfer functions $b(k)$ conditioned on halo environment. It demonstrates that, at fixed halo mass, halos in sheets cluster most strongly. The difference in the clustering of halos residing in filaments and in knots is also significant. These findings are consistent with previous work showing the strong dependence of halo abundances and clustering on $\alpha$ (see Paranjape et al., 2018).
@@
4 Correlation between ELG properties and environment
Having established that the large-scale bias of mock $\mathrm{H}\rm{\alpha}$ emitters strongly correlates with their anisotropy parameter, we now quantify in this section how strongly anisotropy correlates with galaxy properties.
4.1 ELG properties
We will focus on a few galaxy properties (loosely labelled by $I_{g}$) and examine their correlation with the anisotropy parameters $\alpha_{\rm{T}}$ and $\alpha_{\rm{V}}$ evaluated on the host halo scale. In addition to the $\mathrm{H}\rm{\alpha}$ luminosity, we will focus on 5 properties: the instantaneous SFR (star formation rate), the total galactic gas mass $m_{\rm{gas}}$, the 1-dimensional velocity dispersion $\sigma_{V}$ of all particles/cells (i.e. gas, stars, dark matter) bound to the galaxy, the total acccretion rate $\dot{M}_{\rm{BH}}$ of all black holes in the galaxy, and the total galactic gas inflow rate $\dot{m}_{\rm{gas}}$. The first four quantities can be extracted directly from the public galaxy catalogs [^1] SubhaloMassType for the gas mass, SubhaloVelDisp for the velocity dispersion, SubhaloBHMdot for the black hole accretion rate and SubhaloSFR for the total SFR. whereas, for the instantaneous gas inflow rate, we use the expression adopted by Nelson et al. (2019b). Namely, we determine the rate of change in the total galactic gas mass from the Lagrangian properties of gas cells at a given time according to
@@ 4.1 ELG properties
[…]
4.2 Correlation with environment
@@
We carry out a correlation analysis to identify the galaxy properties that correlates most with the anisotropy parameter. Instead of performing a Pearson test (which assumes a linear correlation), we use again the Spearman rank order correlation coefficient $\rho_{s}$ to test for any monotonic relation. Concretely, we compute $\rho_{s}(\alpha\leftrightarrow I_{g})$, where $I_{g}$ is one of the galaxy properties outlined above. A positive $\rho_{s}(\alpha\leftrightarrow I_{g})$ implies that $I_{g}$ tends to increase with $\alpha$ or, according to the top panel of Fig. 6, as one moves from knots to filaments and sheets. Notice that the null probability decreases with increasing value of $|\rho_{s}|$ and sample size $N_{s}$.
[…]
@@
Fig. 8 is identical to Fig. 7 except that $\rho_{s}$ is displayed as a function of the total $\mathrm{H}\rm{\alpha}$ luminosity of the mock ELGs. It is interesting that the trends, when shown as a function of $L_{\mathrm{H}\rm{\alpha}}$, can be very different from those shown as a function of $M$. Consider $L_{\mathrm{H}\rm{\alpha}}^{\text{RR}}$ for instance. In Fig. 7, the corresponding $\rho_{s}$ can be as large as 0.3 for $M\sim 10^{11}\ M_{\odot}/h$ whereas, in Fig. 8, it hardly deviates from zero, regardless of the value of $L_{\mathrm{H}\rm{\alpha}}$. By contrast, $\rho_{s}(\alpha\leftrightarrow L_{\mathrm{H}\rm{\alpha}}^{\text{AGN}})$ for the central galaxies shows a behavior consistent with an approximately constant $M/L_{\mathrm{H}\rm{\alpha}}$ ratio.
@@
Figure 9: Correlations between $\alpha$ and a range of galaxy properties defined in 4.1: the total gas mass, 1-D velocity dispersion (all galaxy member particles), the BH accretion rate, the total stellar mass and the total star formation rate. The solid lines show correlations with $\alpha_{\rm{V}}$ while the dashed show $\alpha_{\rm{T}}$. The top panel shows correlations for central galaxies, while the bottom panel shows for satellite galaxies
@@
Likewise, Fig. 9 displays the Spearman rank order correlation coefficient $\rho_{s}(\alpha\leftrightarrow I_{g})$ for a few galaxy properties $I_{g}=(m_{\rm gas},\sigma_{V},\dot{M}_{\rm{BH}},{\rm SFR})$ extracted directly from the IllustrisTNG simulation data (i.e. they are not post-processed quantities). Here again, the correlations are weak ($|\rho_{s}|$ rarely exceeds 0.2), albeit statistically significant up to $M\sim 10^{12}\ M_{\odot}/h$ at least. The negative $\rho_{s}$ between $\alpha$ and $\sigma_{V}$ presumably reflects the fact that, since $\sigma_{V}$ measures the depth of the potential well, it will be larger for the deeper and more spherically symmetric potential wells of (sub)halos located at the nodes of the cosmic web. Since the latter preferentially have low values of $\alpha$ (as illustrated by the top panel of Fig. 6), the resulting $\rho_{s}$ is negative. By contrast, the positive $\rho_{s}(\alpha\leftrightarrow{\rm SFR})$ (and $\rho_{s}(\alpha\leftrightarrow m_{\text{gas}})$ to a lesser extent) for central galaxies reflects the fact that the total gas content and star formation rate are (slightly) higher in environments with larger anisotropy.
@@
Figure 10: Correlation between $\alpha$ and the gas accretion rate onto individual galaxies for the 3 different settings defined in 4.1. The solid lines show correlations with $\alpha_{\rm{V}}$ while dashed lines show the correlations with $\alpha_{\rm{T}}$. The top panel the shows correlation for central galaxies, while the bottom panel shows for satellite galaxies.
@@
To understand this further, Fig. 10 shows the correlation coefficient $\rho_{s}(\alpha\leftrightarrow\dot{m}_{\text{gas}}^{(i)})$ between $\alpha$ and the gas inflow rate computed at 3 different radius (see §4.1 for details). The correlations are significant for central galaxies solely. In this case, the negative values of $\rho_{s}$ imply that central galaxies with a higher net gas inflow (i.e. a positive $\dot{m}_{\text{gas}}$) are preferentially found in isotropic environments (i.e. low value of $\alpha$), presumably because in highly anisotropic environments accretion is restricted to fewer directions. It is also possible that anisotropic environments contain more angular momentum, which inhibits accretion (e.g. Song et al., 2021). The signal is strongest when $\dot{m}_{\text{gas}}$ is computed at the intermediate radius $4r_{1/2}$. These negative values of $\rho_{s}$ are, however, not at odds with the positive correlation $\alpha\leftrightarrow{\rm SFR}$ reported in Fig. 9 because cooling, turbulence and mixing, cloud condensation etc. generally introduce a delay between gas accretion and star formation.
@@
4.3 Scale-dependent bias
@@
To see this, it is instructive to look at the scale-dependence of $b(k)$ as a function of galaxy properties $I_{g}$. Fig. 12 displays the ratio $b(k)/b_{1}$ for central (solid curves) and satellite (dashed curves) ELGs selected by $L_{\mathrm{H}\rm{\alpha}}^{\text{RR}}$, ${\rm SFR}_{\text{gal}}$ and $\mathrm{H}\rm{\alpha}$ line FWHM (from top to bottom). The mocks have been split into terciles according to the color code shown in the figure. While the satellite galaxies exhibit a strong scale-dependent bias above $k\sim 0.1\ {\rm{\it h}Mpc^{-1}}$, $b(k)$ is weaker for central galaxies.Furthermore, central galaxies in the highest tercile of $L_{\mathrm{H}\rm{\alpha}}^{\text{RR}}$, ${\rm SFR}_{\text{gal}}$ and $\mathrm{H}\rm{\alpha}$ line FWHM have the weakest scale-dependence, while it is the opposite for satellite galaxies. This scale-dependence reflects the fact that bias parameters beyond the linear bias $b_{1}$ also correlate with the anisotropy parameter (see e.g. Ramakrishnan & Paranjape, 2020, for measurements of $b_{2}$ as a function of $\alpha$). Note that, as one transitions from pair counts in separate halos to pairs in the same halo, one expects to see a strong scale-dependent bias. This transition is typically associated with the halo virial radius, for which the corresponding scale is $k\propto 1/r_{\rm vir}$. This is larger than the $k\sim 0.1\ {\rm{\it h}Mpc^{-1}}$ shown here and, therefore, unlikely to explain our satellite measurements — unless the relevant scale is the one that sets $\alpha$.
5 Conclusions

### v2

@@ III.3 Cosmic web classification
The right panel of Fig. 5 displays the environmental mark fractions of ELG parent halos as a function of the ELG total $\mathrm{H}\rm{\alpha}$ luminosity. In each luminosity bin, we computed the fraction of galaxies in every classification, such that the fractions sum to 1 by definition. The threshold values $\lambda_{\text{th,V}}$ and $\lambda_{\text{th,T}}$ adopted here ensure that the V- and T-web classifications are in reasonably good agreement with those predicted by DisPerSE for $\mathrm{H}\rm{\alpha}$ luminosities greater than $10^{41}\ {\rm erg\,s^{-1}}$. Observe also that the fraction of ELGs in filaments is nearly constant and equal to $\approx 40\%$ across four orders of magnitude in luminosity, unlike the fraction of galaxies in sheets. At $L_{\mathrm{H}\rm{\alpha}}>10^{43}\ {\rm erg\,s^{-1}}$, nearly all galaxies are in knots. Finally, note that ELGs with parent halo mass smaller than $2\times 10^{11}\ {\rm{\it h}}^{-1}M_{\odot}$ are not included in this classification due to the limited resolution of the grid. This excludes only $\approx 18\%$ ($2\%$) of ELGs with $L_{\mathrm{H}\rm{\alpha}}>10^{39}\ {\rm erg\,s^{-1}}$ ($L_{\mathrm{H}\rm{\alpha}}>10^{42}\ {\rm erg\,s^{-1}}$).
III.4 Environmental bias
@@
The peak-background split argument relates the magnitude of the large scale bias $b_{1}$ to (minus) the slope of the halo mass function (Kaiser, 1984; Bardeen et al., 1986; Cole and Kaiser, 1989; Mo and White, 1996; Sheth and Tormen, 1999). In light of this, the left panel of Fig. 5 suggests that, at fixed halo mass, halos in sheets and filaments should be significantly more clustered than those residing in knots.
@@
To illustrate this point, the top panel of Fig. 6 shows the distribution of the anisotropy parameters $\alpha_{\rm V}$ and $\alpha_{\rm T}$ for halos with mass $M\geq 5\times 10^{11}\ M_{\odot}/h$. The anisotropy parameter thus is a good discriminator of the environment. Furthermore, the largest $\alpha_{\rm V}$ and $\alpha_{\rm T}$ values are exclusively associated with the rare halos residing in sheets. The bottom panel of Fig. 6 displays the corresponding bias transfer functions $b(k)$ conditioned on halo environment. It demonstrates that at fixed halo mass, halos in sheets cluster most strongly, even though halos in knots and filaments tend to be more massive on average. In the T-web classification, the mean halo masses are $\approx 10^{10.4}$, $10^{10.5}$, $10^{11.1}$ and $10^{11.6}\ {\rm{\it h}}^{-1}M_{\odot}$ for halos in voids, sheets, filaments and knots, respectively. The difference in the clustering of halos residing in filaments and in knots is also significant. These findings are consistent with previous work showing the strong dependence of halo abundances and clustering on $\alpha$ (see Paranjape et al., 2018).
@@
IV Correlation between ELG properties and environment
Having established that the large-scale bias of ELGs strongly correlates with their anisotropy parameter, we now quantify in this section how strongly anisotropy correlates with galaxy properties.
IV.1 ELG properties
We will focus on a few galaxy properties (loosely labelled by $I_{g}$) and examine their correlation with the anisotropy parameters $\alpha_{\rm{T}}$ and $\alpha_{\rm{V}}$ evaluated on the host halo scale. In addition to the $\mathrm{H}\rm{\alpha}$ luminosity, we will focus on 5 properties: the instantaneous SFR (star formation rate), the total galactic gas mass $m_{\rm{gas}}$, the 1-dimensional velocity dispersion $\sigma_{V}$ of all particles/cells (i.e. gas, stars, dark matter) bound to the galaxy, the total acccretion rate $\dot{M}_{\rm{BH}}$ of all black holes in the galaxy, and the total galactic gas inflow rate $\dot{m}_{\rm{gas}}$. The first four quantities can be extracted directly from the public galaxy catalogs [^1] SubhaloMassType for the gas mass, SubhaloVelDisp for the velocity dispersion, SubhaloBHMdot for the black hole accretion rate and SubhaloSFR for the total SFR. whereas, for the instantaneous gas inflow rate, we use the expression adopted by Nelson et al. (2019a). Namely, we determine the rate of change in the total galactic gas mass from the Lagrangian properties of gas cells at a given time according to
@@ IV.1 ELG properties
which, up to sign, is a discretized form of the expression $\dot{m}=\int_{S}\rho(\vec{v}\cdot\hat{n})dS$ that encodes conservation of mass, where $S$ is some 2-dimensional surface with perpendicular unit vector $\hat{n}$; $\rho,\vec{v}$ are the gas density and velocity respectively. Here, the sum runs over all the gas cells bound to the galaxy, which lie in a spherical shell of radius $r_{0}$ and width $\Delta r$ centered on the galaxy position (defined by the position of the most strongly bounded particle). Furthermore, $\vec{r}$ and $\vec{v}$ are the physical position and velocity of the particles (relative to the galaxy frame), $r=|\vec{r}|$, and $m_{\text{gas,cell}}$ is the gas cell mass. Therefore, a positive value of $\dot{m}_{\text{gas}}$ indicates a net inflow of gas inside the shell of radius $r_{0}$. Note that we scale $r_{0}$ with the galaxy (total) half-mass radius $r_{1/2}$ to account for variations in the size of galaxies. For each galaxy, we compute three gas inflow rates $\dot{m}_{\text{gas}}^{(1)}$, $\dot{m}_{\text{gas}}^{(2)}$ and $\dot{m}_{\text{gas}}^{(3)}$ on assuming $r_{0}=2r_{1/2}$, $4r_{1/2}$ and $7r_{1/2}$, respectively. The width is scaled according to $\Delta r=r_{0}/4$.
IV.2 Correlation with environment
@@
We carry out a correlation analysis to identify the galaxy properties that correlate most with the anisotropy parameter. Instead of performing a Pearson test (which assumes a linear correlation), we use again the Spearman rank order correlation coefficient $\rho_{s}$ to test for any monotonic relation. Concretely, we compute $\rho_{s}(\alpha\leftrightarrow I_{g})$, where $I_{g}$ is one of the galaxy properties outlined above. A positive $\rho_{s}(\alpha\leftrightarrow I_{g})$ implies that $I_{g}$ tends to increase with $\alpha$ or, according to the top panel of Fig. 6, as one moves from knots to filaments and sheets. Notice that the null probability decreases with increasing value of $|\rho_{s}|$ and sample size $N_{s}$.
The left panel of Fig. 7 shows the correlation coefficient $\rho_{s}$ between the anisotropy parameter and the different sources of $L_{\mathrm{H}\rm{\alpha}}$ luminosity in our emission line model. The solid and dashed curves were obtained with $\alpha_{\rm V}$ and $\alpha_{\rm T}$, respectively. They are shown as functions of halo mass for central (top panel) and satellite galaxies (bottom panel). The shaded areas show the uncertainty $\delta\rho_{s}$ on the measurement of $\rho_{s}$. It is computed according to the formula $\delta\rho_{s}=\sqrt{\frac{1-\rho_{s}^{2}}{N_{s}-2}}$. To avoid clutter, error bars are shown for only one source of $\mathrm{H}\rm{\alpha}$ emission.
The correlations weaken as $M$ increases significantly above $M_{*}$. For $M\sim M_{\star}$, $\alpha$ correlates most strongly with $L_{\mathrm{H}\rm{\alpha}}^{\rm RR}$ for central galaxies, whereas for satellite galaxies, the strongest correlation is with $L_{\mathrm{H}\rm{\alpha}}^{\rm AGN}$. The correlation of $\alpha$ and $L_{\mathrm{H}\rm{\alpha}}^{\rm AGN}$ indicates that satellite galaxies are influenced by AGN activity in nearby centrals (AGN are mostly found in central galaxies). This influence appears to suppress star formation in small satellite ELGs more efficiently in large $\alpha$ environments.
The correlation between $\alpha$ and $L_{\mathrm{H}\alpha}^{\mathrm{RR}}$ is consistent with a physical picture in which galaxies embedded in sheets and filaments experience large-scale accretion shocks. Such shocks are theoretically expected (Mo et al., 2010) and are observed in cosmological simulations (Mandelker et al., 2019; Ramsøy et al., 2021; Pasha et al., 2023). They can influence galaxy evolution by heating and ionizing the circumgalactic gas and thus quenching star formation (Hasan et al., 2023, 2024). In the TNG galaxy formation model, ELGs residing in low-mass halos within these environments accrete gas from a hot, ionized phase that produces $\mathrm{H}\alpha$ emission through radiative recombinations as it cools and condenses into the ISM. Therefore, we interpret the $\alpha\leftrightarrow L_{\mathrm{H}\alpha}^{\mathrm{RR}}$ trend as a manifestation of this environment-driven accretion regime, with the caveat that this is a simulation-informed scenario rather than a model-independent prediction.
The right panel of Fig. 7 shows the correlation between $\alpha$ and the $\mathrm{H}\rm{\alpha}$ line full width at half maximum (FWHM; see §A for details of its computation) for central (top panel) and satellite galaxies (bottom panel). For centrals, a weak but non-vanishing correlation is detected up to $M\sim 10^{12}\ M_{\odot}/h$. In Appendix §B, we have checked the numerical convergence of several of these correlations and found them to be robust down to a mass $M\sim{\rm a~few}\ 10^{10}\ {\rm{\it h}}^{-1}M_{\odot}$.
@@
Fig. 8 is identical to Fig. 7 except that $\rho_{s}$ is displayed as a function of the total $\mathrm{H}\rm{\alpha}$ luminosity of the mock ELGs. It is interesting that the trends, when shown as a function of $L_{\mathrm{H}\rm{\alpha}}$, can be very different from those shown as a function of $M$. Consider $L_{\mathrm{H}\rm{\alpha}}^{\text{RR}}$ for instance. In Fig. 7, the corresponding $\rho_{s}$ can be as large as 0.3 for $M\sim 10^{11}\ M_{\odot}/h$ whereas, in Fig. 8, $|\rho_{s}|$ does not exceed 0.1, regardless of the value of $L_{\mathrm{H}\rm{\alpha}}$. By contrast, plotting $\rho_{s}(\alpha\leftrightarrow L_{\mathrm{H}\rm{\alpha}}^{\text{PE}})$ as a function of $L_{\mathrm{H}\rm{\alpha}}$ reveals a negative correlation at low luminosities where PE dominates (Rapoport et al., 2025). This trend is consistent with the physical picture outlined above: star formation is quenched in filamentary environments, which suppresses the PE rate.
@@
Figure 9: Correlations between $\alpha$ and a range of galaxy properties defined in IV.1: the total gas mass, 1-D velocity dispersion (all galaxy member particles), the BH accretion rate, the total stellar mass and the total star formation rate. The solid lines show correlations with $\alpha_{\rm{V}}$ while the dashed show $\alpha_{\rm{T}}$. The top panel shows correlations for central galaxies, while the bottom panel shows for satellite galaxies
@@
Likewise, Fig. 9 displays the Spearman rank order correlation coefficient $\rho_{s}(\alpha\leftrightarrow I_{g})$ for a few galaxy properties $I_{g}=(m_{\rm gas},\sigma_{V},\dot{M}_{\rm{BH}},{\rm SFR})$ extracted directly from the IllustrisTNG simulation data (i.e. they are not post-processed quantities). Here again, the correlations are weak ($|\rho_{s}|$ rarely exceeds 0.2), albeit statistically significant up to $M\sim 10^{12}\ M_{\odot}/h$ at least. The negative $\rho_{s}$ between $\alpha$ and $\sigma_{V}$ reflects the fact that $\sigma_{V}$ measures the depth of the potential well: (sub)halos located at the knots of the cosmic web have deeper, more spherically symmetric potential wells and, therefore, larger $\sigma_{V}$. As shown in the top panel of Fig. 6, these environments preferentially host halos with low $\alpha$, leading to the observed negative correlation. By contrast, the positive $\rho_{s}(\alpha\leftrightarrow{\rm SFR})$ (and $\rho_{s}(\alpha\leftrightarrow m_{\text{gas}})$ to a lesser extent) for central galaxies reflects the fact that the total gas content and star formation rate are (slightly) higher in environments with larger anisotropy. Finally, the positive $\rho_{s}(\alpha\leftrightarrow\dot{M}_{\rm{BH}})$ at halo mass $M\lesssim 10^{11}\ {\rm{\it h}}^{-1}M_{\odot}$ explains the (weak) trend between $\alpha$ and $L_{\mathrm{H}\alpha}^{\mathrm{AGN}}$ seen in the bottom panel of Fig. 7. A larger $\alpha$ leads to a larger $\dot{M}_{\rm{BH}}$ which, in turn, increases the AGN flux incident on nearby satellite galaxies. This enhances their photo-excitation rate and thus raises $L_{\mathrm{H}\alpha}^{\mathrm{AGN}}$.
@@
Figure 10: Correlation between $\alpha$ and the gas accretion rate onto individual galaxies for the 3 different settings defined in IV.1. The solid lines show correlations with $\alpha_{\rm{V}}$ while dashed lines show the correlations with $\alpha_{\rm{T}}$. The top panel the shows correlation for central galaxies, while the bottom panel shows for satellite galaxies.
@@
To understand this further, Fig. 10 shows the correlation coefficient $\rho_{s}(\alpha\leftrightarrow\dot{m}_{\text{gas}}^{(i)})$ between $\alpha$ and the gas inflow rate computed at 3 different radius (see §IV.1 for details). The correlations are significant for central galaxies solely. In this case, the negative values of $\rho_{s}$ imply that central galaxies with a higher net gas inflow (i.e. a positive $\dot{m}_{\text{gas}}$) are preferentially found in isotropic environments (i.e. low value of $\alpha$) because, in highly anisotropic environments, accretion is restricted to fewer directions. It is also possible that anisotropic environments contain more angular momentum, which inhibits accretion (e.g. Song et al., 2021). The signal is strongest when $\dot{m}_{\text{gas}}$ is computed at the intermediate radius $4r_{1/2}$. These negative values of $\rho_{s}$ are, however, not at odds with the positive correlation $\alpha\leftrightarrow{\rm SFR}$ reported in Fig. 9 because cooling, turbulence and mixing, cloud condensation etc. generally introduce a delay between gas accretion and star formation.
@@
IV.3 Scale-dependent bias
@@
To see this, it is instructive to look at the scale-dependence of $b(k)$ as a function of galaxy properties $I_{g}$. Fig. 12 displays the ratio $b(k)/b_{1}$ for central (solid curves) and satellite (dashed curves) ELGs selected by $L_{\mathrm{H}\rm{\alpha}}^{\text{RR}}$, ${\rm SFR}_{\text{gal}}$ and $\mathrm{H}\rm{\alpha}$ line FWHM (from top to bottom). The mocks have been split into terciles according to the color code shown in the figure. While the satellite galaxies exhibit a strong scale-dependent bias above $k\sim 0.1\ {\rm{\it h}Mpc^{-1}}$, $b(k)$ is weaker for central galaxies. Furthermore, central galaxies in the highest tercile of $L_{\mathrm{H}\rm{\alpha}}^{\text{RR}}$, ${\rm SFR}_{\text{gal}}$ and $\mathrm{H}\rm{\alpha}$ line FWHM have the weakest scale-dependence, while it is the opposite for satellite galaxies. This scale-dependence reflects the fact that bias parameters beyond the linear bias $b_{1}$ also correlate with the anisotropy parameter (see e.g. Ramakrishnan and Paranjape, 2020, for measurements of $b_{2}$ as a function of $\alpha$). Note that, as one transitions from pair counts in separate halos to pairs in the same halo, one expects to see a strong scale-dependent bias. This transition is typically associated with the halo virial radius, for which the corresponding scale is $k\propto 1/r_{\rm vir}$. This is larger than the $k\sim 0.1\ {\rm{\it h}Mpc^{-1}}$ shown here and, therefore, unlikely to explain our satellite measurements — unless the relevant scale is the one that sets $\alpha$.
V Conclusions